What is the Fermi Wavelength of 2DEG in an AlGaAs/GaAs hetero-coupled HEMT substrate with a low-temperature mobility of 100 m^2/Vs and an electron density of 5*10^15/m^2? And what is the mean free time τ of 2DEG when the effective mass of 2DEG is 0.03? And how much is the mean free path of 2DEG?
Answer #1

Glven that. 15 u= 5X10 - 2 m2 n=sxio m = 0.0 0.03 Kg ma m=0.0 mo mois ēmasswe know that с et 100= 1-GX10 19 z 0.03% 9.1x1531 so ī= -29 0.03 x 9 al X10 10681019 ここ sec. 0.170625x1010 I= 17.0625 psecih